OpenStudy (anonymous):

PLEASEE HELP. I NEED HELP FAST..PLEASE ;( Identify the 8th term of the geometric sequence -2, -12, -72 (2 points) -3,359,232 -559,872 559,872 3,359,232 Identify the 12th term of a geometric sequence where a1 = 8 and a6 = -8,192. (2 points) 134,217,728 33,554,432 -33,554,432 -134,217,728

OpenStudy (anonymous):

@mathstudent55 @phi , Can yall please help out. Show me the steps i Really need help fast :( !

OpenStudy (anonymous):

- n * 6 -2*6 =-12 -12*6= -72 -72*6= -432 -432*6= -2592 -2592*6= -15552 -15552*6= -93312 -93312*6= -559872 I don't know the second one

OpenStudy (anonymous):

I forgot to say the answer to the first one is -559872.

OpenStudy (anonymous):

Thank You So Much